OpenStudy (anonymous):

i know 8 cubed is 512 but how did i get that answer?? I just know it

OpenStudy (anonymous):

\[8^{3} = 8x8x8 = 512\]

OpenStudy (anonymous):

calculater

hero (hero):

You multiplied \(8 \times 8 \times 8\)

OpenStudy (anonymous):

8x8x8 = 512 ex: 2x2x2 = 8 = 2^3

OpenStudy (anonymous):

or 2x2x2x2= 2^4

OpenStudy (anonymous):

thanks! the exponents confuse me, I can't use a calculator for this college algebra class.

OpenStudy (anonymous):

cube means multiplying the same no. thrice